What is the molar concentration of carbonate ion in a 0.10 M solution of carbonic acid?

  Ka1 = 4.3 × 10–7 ; Ka2 = 5.6 × 10–11

Please report the answer in units of M (do not write unit, only number)

Write numbers as 1.23E-5
Answer #1

H2CO3 +420 ² HCOğ + H₂ot ka, = 4.3x10 7 HCO3 + H2O È což +4got Rag= 5.6x10- consider first dissoci ation g 4,603 H₂CO3 + H2OS Ka= [co3²] [430f) (मी 5.6X10!! = 1 0.000207 +) (v) 10.000207-2) x is small so neglect 5.6x10-11 = - 0.000204) x 0.000 got x